# Stochastic Geometry and Dynamical System Analysis of Walker Satellite Constellations

### v2

@@ V Conclusion
In this paper, we developed a stochastic geometry model for the well-known Walker satellite constellation. We first showed its time invariance and then proved other essential statistical properties of the dynamical system, such as conditions for its periodicity and ergodicity. We then derived the typical performance of downlink communications as a function of key parameters, including the number of orbits $N_{o}$, the number of satellites $N_{s}$, the orbit inclination $\phi$, and user latitude $l_{u}$. The derived metrics were interpreted based on the properties obtained from the dynamical system analysis. The proposed framework not only provides a tractable way to understand the interplay between network variables but also offers an intuitive method for designing and optimizing LEO satellite networks, grounded in the dynamical aspects of such systems.
@@
The established framework can be extended to incorporate multi-class Walker constellations, which are used in real-world forthcoming satellite networks. Future research may explore a variety of applications where satellites follow Walker-type configurations. For example, the framework could be applied to analyze data harvesting via satellite relays or to investigate and optimize latitude-dependent Internet routing.

### v3

@@ V Conclusion
In this paper, we established a stochastic geometry model for the Walker satellite constellation. We identified time invariance and proved other structural properties of the dynamical system, such as its periodicity and ergodicity. We then derived the typical performance of downlink communications as a function of key parameters, including the number of orbits, the number of satellites, the orbit inclination, and the user latitude. The derived metrics were interpreted based on the properties obtained from the dynamical system analysis. The proposed framework not only provides a tractable way to understand the interplay between network variables but also offers an intuitive method for designing and optimizing LEO satellite networks. Note the network performance with nonzero phase difference can be analyzed using similar techniques.
@@
The established framework can be used to explore various applications based on multi-class Walker constellations. The framework could be used to analyze data harvesting via satellite relays or to investigate Internet routing.
Appendix A No Satellite Collision
Lemma 1.
Consider the Walker constellation with parameters $\phi\neq 0$, $N_{0}$, and $N_{s}$.
  • Consider first antipodal pairs of orbits (if the number of orbits $N_{o}$ is even, then there are $N_{o}/2$ such pairs; if it is odd, then there are none). If the number of satellites $N_{s}$ is even as well, then the two antipodal satellites of any two antipodal orbits collide. If $N_{s}$ is odd, then there is no collision on antipodal orbits.
  • Consider now non-antipodal pairs of orbits. Each such pair is characterized by an integer $k$, with $0<k<\frac{N_{o}}{2}$, where $\frac{k2\pi}{N_{o}}$ is the angle (on the equator) between the two descending and ascending points of the orbits of such a pair. Then, a pair of satellites on these two orbital planes collide if and only if $\tan\left(\frac{l\pi}{N_{s}}\right)=\cos(\phi)\tan\left(\frac{k\pi}{N_{o}}\right)$ for some integer $0<l<\frac{N_{s}}{2}$.
Proof:
Orbits are numbered from 0 to $N_{0}-1$. The orbit 0 is the reference orbit. Let $0<k\leq N_{o}-1$ be another orbit. Note that if there are no collisions between orbit 0 and orbit $0<k\leq\frac{N_{o}}{2}$, then there are no collisions between orbit $\frac{N_{o}}{2}<k\leq N_{o}-1$ and orbit 0, by the rotational invariance. Hence it is enough to check collisions between orbit 0 and orbit $k$, with $0<k\leq\frac{N_{o}}{2}$.
Let $A_{k}$ (resp $B_{k}$) be the ascending (resp. descending) point of orbit $k$, with $0<k\leq\frac{N_{o}}{2}$.
We first consider the case where $N_{0}$ is even and $k=N_{o}/2$. In this case, the two orbits are antipodal and they cross the equator at the same points. Assume that at some given time, there is a satellite of the orbit 0 at $A_{0}$. Then, either $N_{s}$ is even too and there is then a satellite of orbit $k$ at $B_{k}$. Since $A_{0}=B_{k}$, there is hence a collision between satellites of orbit 0 and of orbit $N_{0}/2$, regardless of the inclination $\phi$. If $N_{s}$ is odd, there is then no collision on these two orbits. This proves the first statement.
Consider now two orbits 0 and $k$, with $0<k<\frac{N_{o}}{2}$ (this is equilalent to all $0<k\leq\frac{N_{o}}{2}$ if $N_{o}$ is even and all $0<k\leq\frac{N_{o}}{2}$ other than the antipodal orbits otherwise). These two orbits meet at two points. Let $C_{k}$ be the intersection point on the northern hemisphere. Consider the spherical triangle $\triangle A_{k}C_{k}B_{0}$, Then, the longitudinal difference between $A_{k}$ and $B_{0}$ is given by $c=\pi-\frac{2\pi}{N_{o}}k$.
Let $a$ be the length of the arc ¿ $A_{k}C_{k}$ Then, based on the definition of the inclination, we have $\sphericalangle C_{k}A_{k}B_{0}=\sphericalangle C_{k}B_{0}A_{k}=\phi$, namely $\triangle A_{k}C_{k}B_{0}$ is a isosceles spherical triangle. Hence the length of the arc ¿ $C_{k}B_{0}$ is equivalent to $a$. Moreover, since $a<\pi$, we can use Napier’s analogy [15] for a isosceles spherical triangle $\alpha=\beta=\phi$ to get
$\cot(a)$ $=\cos(\phi)\frac{1+\cos(c)}{\sin(c)}$
$=\cos(\phi)\cot(c/2)$
$=\cos(\phi)\cot\left(\frac{\pi}{2}-\frac{k\pi}{N_{o}}\right)$
$=\cos(\phi)\tan\left(\frac{k\pi}{N_{o}}\right).$ (18)
Then, there is a collision at $C_{k}$ between satellites of these two orbits if and only if there is an integer $l$ such that
$\pi-2a=\frac{2\pi}{N_{s}}l,\quad\mbox{mod }2\pi.$ (19)
[…]
As a result, for $0<a<\pi,$ there is a collision if and only if for some $l$,
$2a=\pi-\frac{l2\pi}{N_{s}}\quad\mbox{mod }2\pi$ (20)
or equivalenly,
$a=\frac{\pi}{2}-\frac{l\pi}{N_{s}}\quad\mbox{mod }\pi.$ (21)
Since $0<a<\pi$, we have $-\frac{N_{s}}{2}<l<\frac{N_{s}}{2}$.
Hence, using Eq. (A), we conclude that there is a collision if and only if there exsists an integer $k$ with $0<k<\frac{N_{o}}{2}$ and an integer $l$ with $-\frac{N_{s}}{2}<l<\frac{N_{s}}{2}$ such that
$\tan\left(\frac{l\pi}{N_{s}}\right)=\cos(\phi)\tan\left(\frac{k\pi}{N_{o}}\right),$
Note that since $\cos(\phi)\geq 0$, if such an $l$ exists, it must be positive. ∎
A direct practical implication of the above lemma is as follows: for all integers $N_{s}$ and $N_{o}$, if one of them is odd, one can then always find an orbit inclination $\phi^{\star}$ arbitrarily close to a given $\phi$, and such that there is no collision of satellites at any time in the proposed Walker constellation model.